# Beyond the Linear Separability Ceiling

### v1

@@ Appendix I Results across HOI splits
LoRA ($\mathcal{L}_{\text{NT}}$) $69.0\pm 3.2$ $74.1\pm 3.0^{G}$ $50.0$
LoRA ($\mathcal{L}_{\text{combined}}$) $65.8\pm 3.3$ $74.1\pm 3.0^{G}$ $67.4\pm 3.2^{G}$
Across all models, a key observation in cross-domain generalization is that the generative accuracy remains tightly coupled with the final representation classification accuracy when trained with the combined loss (($\mathcal{L}_{\text{combined}}$).
Conditional generalization of Phi. It successfully surpasses its LSC when trained on the broader HOI dataset and tested on OpenWorld, but this success is not bidirectional.
Robust generalization of Pixtral. It consistently surpasses its LSC when transferring between domains when trained with $\mathcal{L}_{\text{combined}}$, succeeding via representation refinement. This ability to significantly improve performance in both transfer directions points to a more fundamentally sound approach for generalizable reasoning.
Appendix K Class imbalance investigation
Evaluating with unbalanced classes can lead to a misleading estimate of the LSC. This can be most easily seen when we split the accuracy into two, positive and negative only ground labels, as done in Table 10.
Table 10: Class imbalances across datasets and models
Model Dataset Method
gen. acc.
pos. (%)
gen. acc.
neg. (%)
gen. acc.
avg. (%)
LSC
(pos., %)
LSC
(neg., %)
LSC
(avg., %)
Phi Openworld Direct (interleaved) 75.2 42.8 59.0 86.8 81.2 84.0
LoRA ($\mathcal{L}_{\text{NT}}$) 91.2 88.4 89.8 86.8 81.6 84.2
LoRA ($\mathcal{L}_{\text{combined}}$) 94.4 98.0 96.2 87.2 81.2 84.2
Hoi Direct (interleaved) 97.5 6.8 52.1 75.0 68.8 71.9
LoRA ($\mathcal{L}_{\text{NT}}$) 76.2 78.5 77.4 74.2 68.0 71.1
LoRA ($\mathcal{L}_{\text{combined}}$) 82.0 75.8 78.9 75.2 68.2 71.8
Pixtral Openworld Direct (interleaved) 80.8 64.0 72.4 78.4 73.6 76.0
LoRA ($\mathcal{L}_{\text{NT}}$) 92.8 94.0 93.4 79.2 74.0 76.6
LoRA ($\mathcal{L}_{\text{combined}}$) 95.6 94.4 95.0 76.4 72.4 74.4
Hoi Direct (interleaved) 67.8 47.8 57.8 64.8 60.8 62.7
LoRA ($\mathcal{L}_{\text{NT}}$) 80.5 75.5 78.0 62.5 60.8 61.6
LoRA ($\mathcal{L}_{\text{combined}}$) 80.5 78.8 79.6 67.5 58.8 63.1
Gemma3 4B Openworld Direct (interleaved) 69.6 82.4 76.0 92.4 87.2 89.8
LoRA ($\mathcal{L}_{\text{NT}}$) 92.0 92.8 92.4 92.4 87.2 89.8
LoRA ($\mathcal{L}_{\text{combined}}$) 95.6 95.6 95.6 92.4 87.2 89.8
Hoi Direct (interleaved) 24.5 88.5 56.5 77.5 70.8 74.1
LoRA ($\mathcal{L}_{\text{NT}}$) 86.5 82.0 84.2 77.5 70.8 74.1
LoRA ($\mathcal{L}_{\text{combined}}$) 81.5 87.0 84.2 77.5 70.8 74.1
The data in Table 10 demonstrates how average accuracies, whether generative or similarity-based, can mask significant performance disparities between classes. An aggregated LSC is composed of differing per-class separability scores (e.g., for Gemma3 on HOI, 77.5% for positive vs. 70.8% for negative). This class-dependent variability means the overall LSC may not fully capture the model’s representational quality for all distinct classes. Therefore, this detailed per-class separability analysis serves as a valuable diagnostic tool. It effectively uncovers such representational biases that overall average metrics might otherwise conceal.
Appendix L Example semantic concepts
@@
Appendix M Visualizing the Mechanism of PEFT via Attention Maps
@@
M.1 Distinct architectural signatures and tuning mechanisms
@@
Gemma3: overcoming representational degradation
As established in the main paper, Gemma3 models exhibit a high LSC but suffer from severe representational degradation, where the final-layer embeddings ($acc_{final}$) become poorly separated. The attention maps in Figure 5 reveal the architectural reason and the solution. Gemma3 uses an efficient sliding window attention for the most part (narrow diagonal band) (layer 24 behaved differently for some reason).
@@
    With $\mathcal{L}_{\text{combined}}$ tuning, the contrastive loss induces a change that makes the model’s reasoning process visible. In the final layers, the attention map develops more structured, bright vertical stripes. Each well-defined stripe represents a targeted “read” operation, where the model globally accesses the compressed representation of visual features. This structured, cross-image comparison provides a clear mechanistic explanation for how the model overcomes its baseline shortcoming. The explicit contrastive pressure forces it to develop and execute this comparison strategy, repairing its degraded final representations and aligning with the recovery of $acc_{final}$ seen in Table 3.
@@
Figure 5: Attention maps for Gemma3 4B. The contrastive loss in (b) visibly forces global cross-image attention in the final layers, a pattern not as pronounced in (a). This visualizes the mechanism that improves the model’s internal representations.
Phi and Pixtral: representation refinement and feature enhancement
Phi and Pixtral employ a standard causal attention mask across all layers (Figures 6 and 7). Their path to success is not about fixing a shortcoming but refining existing processes.
Phi possessing a higher-quality LSC from the outset, Phi’s fine-tuning is primarily a process of refinement and noise reduction. The attention maps visualize this as the patterns within each image block becoming more structured and less diffuse. Instead of a scattered focus across all visual patches, the model learns to consistently attend to the most salient features while ignoring irrelevant ones. This refined intra-image feature extraction allows for a cleaner signal to be passed to the final layers, enabling the model to better leverage its already-strong representations and surpass its LSC.
@@
Figure 6: Attention maps for the Phi model after LoRA tuning with (a) the $\mathcal{L}_{\text{NT}}$ objective and (b) the $\mathcal{L}_{\text{combined}}$ objective. Both training methods result in a similar outcome: a refinement of attention over the model’s already high-quality initial representations. The patterns become visibly more structured and less diffuse, visualizing a mechanism of noise reduction and more targeted feature aggregation that enables the model to surpass its LSC.
Pixtral’s success, even at baseline, stems from its inherent ability to perform representation refinement. The attention maps reveal the mechanism: an intensified focus on intra-image processing in later layers, visible as more structured diagonal blocks. Crucially, this visual signature is nearly identical when fine-tuning with either the standard $\mathcal{L}_{\text{NT}}$ or the $\mathcal{L}_{\text{combined}}$ objective. This equivalence demonstrates that standard next-token prediction is sufficient to fully engage Pixtral’s innate pathway for extracting and enhancing salient features.
@@
Figure 7: Attention maps for the Pixtral model after LoRA tuning with (a) the $\mathcal{L}_{\text{NT}}$ objective and (b) the $\mathcal{L}_{\text{combined}}$ objective. The attention patterns are nearly identical, showing that both objectives activate the same underlying mechanism.

### v3

@@ Appendix I Combined loss hyperparameter search
LoRA (OpenWorld, $\mathcal{L}_{\text{NT}}$) $69.0\pm 3.2$ $74.1\pm 3.0^{G}$ $50.0$
LoRA (OpenWorld, $\mathcal{L}_{\text{combined}}$) $65.8\pm 3.3$ $74.1\pm 3.0^{G}$ $67.4\pm 3.2^{G}$
The results indicate that models can generalize their reasoning strategies to novel concepts within the same task paradigm. Pixtral demonstrates the most robust generalization, surpassing its LSC in both transfer directions. Phi shows strong generalization when trained on the broader relational concepts of HOI and tested on OpenWorld, but this success is not bidirectional. Across all models, a key observation is that generative accuracy remains tightly coupled with the final representation classification accuracy when trained with the combined loss ($\mathcal{L}_{\text{combined}}$), suggesting the learned alignment is a general property, not specific to the training data distribution
M.2 Generalization to Winoground
To test if fine-tuning imparts a more fundamental reasoning capability, we evaluated the models on the Winoground benchmark without task-specific adaptation. This benchmark provides a compelling test case, as its text-retrieval component directly evaluates the inter-image comparison skills honed during our training. The results in Table A.8 reveal how the nature of the training task and the fine-tuning objective dictates the success of cross-task generalization.
Table A.8: Zero-shot cross-task generalization to the Winoground dataset. The CLIP baseline is a linear probe on the vision encoder of the Phi model (and its respective text encoder), serving as an LSC-like benchmark for that model.
Model Train dataset Objective Text retrieval acc. (%) Image retrieval acc. (%)
Phi Baseline 16.75 2.50
CLIP Baseline (ViT-L/14) 27.75 11.75
OpenWorld $\mathcal{L}_{\text{NT}}$ 24.00 14.75
$\mathcal{L}_{\text{combined}}$ 18.75 1.75
HOI $\mathcal{L}_{\text{NT}}$ 18.25 13.50
$\mathcal{L}_{\text{combined}}$ 29.25 14.00
Pixtral Baseline 28.75 4.25
OpenWorld $\mathcal{L}_{\text{NT}}$ 32.75 19.75
$\mathcal{L}_{\text{combined}}$ 33.25 8.50
HOI $\mathcal{L}_{\text{NT}}$ 43.50 24.75
$\mathcal{L}_{\text{combined}}$ 54.75 12.25
Gemma3 4B Baseline 5.25 0.75
OpenWorld $\mathcal{L}_{\text{NT}}$ 5.75 1.00
$\mathcal{L}_{\text{combined}}$ 12.25 2.75
HOI $\mathcal{L}_{\text{NT}}$ 5.75 3.25
$\mathcal{L}_{\text{combined}}$ 12.25 13.25
Task-specific skill transfer. Generalization is highly dependent on the alignment between the skills cultivated during training and those required by the evaluation task. Fine-tuning on OpenWorld, which focuses on atomic semantic concepts, resulted in poor transfer to the compositional reasoning required by Winoground. This mirrors our findings in the cross-Bongard evaluation, suggesting that the model learned a task-specific strategy that was not broadly applicable to compositional challenges.
Objective-driven alignment. Training on the relational HOI dataset proved far more effective for transferring to Winoground. Within this setting, the $\mathcal{L}_{\text{combined}}$ objective consistently improved text retrieval accuracy over the standard $\mathcal{L}_{\text{NT}}$ loss across all models. This is most pronounced for Pixtral, whose accuracy jumps from 43.50% to 54.75%. This demonstrates that our combined objective successfully aligns the model’s structured visual representations and its generative pathway.
Re-emergence of the LSC. The Winoground evaluation of the Phi model offers a compelling illustration of our framework in a new domain. The untuned model’s generative performance (16.75%) exhibits an alignment gap, falling well below the LSC (27.75%) established by a linear probe on its vision and respective text encoder. Fine-tuning on the relational HOI task with our $\mathcal{L}_{\text{combined}}$ objective closes this deficit, elevating the end-to-end performance to 29.25%. However, this improvement is not statistically significant over the LSC. Consequently, by our framework’s definition, an alignment gap persists. This indicates that the model has learned to exploit the linearly decodable information within its representations but has not yet developed a sophisticated (non-linear) reasoning capability that provides a performance advantage beyond this ceiling on a new task. We hypothesize that achieving such an improvement would require fine-tuning on a dataset that more comprehensively covers the specific compositional and relational concepts tested by the Winoground benchmark.
Appendix N Training sensitivity to auxiliary contrastive objective
To demonstrate the impact of loss ($\mathcal{L}_{combined}=w_{n}\mathcal{L}_{NT}+w_{c}\mathcal{L}_{sim}$) counterparts, we conducted a hyperparameter sweep on the Phi-3.5 model using the HOI dataset on a fixed seed. We compare two scheduling strategies: a constant schedule, where $w_{c}$ remains fixed throughout training and $w_{n}$ is fixed to 1, and a dual cosine schedule (detailed in Appendix I), designed to provide a strong initial structural signal that gradually yields to the next-token prediction objective. Results are summarized in Table A.9.
Table A.9: Sensitivity analysis of $\mathcal{L}_{combined}$ hyperparameters on Phi-3.5 (HOI dataset). We compare the constant schedule against dual cosine schedule. The CLIP baseline is a linear probe on the vision encoder of the Phi model (and its respective text encoder), serving as an LSC-like benchmark for that model. Final rep. denotes the linear separability of final representations, and Wino. denotes zero-shot text and image retrieval accuracies on Winoground. The in-distribution (ID) test uses interleaved prompt structure seen during training, while the out-of-distribution (OOD) test uses a labeled prompt structure.
Method Schedule Generative Generative Final rep. Final rep. Wino. text Wino. image
(ID, %) (OOD, %) (ID, %) (OOD, %) (%) (%)
Direct baseline 52.1 58.8 60.5 60.6 16.75 2.50
CLIP baseline (ViT-L/14) 71.9 71.9 27.75 11.75
$w_{c}=0.0$ ($\mathcal{L}_{NT}$) constant 76.00 68.63 61.63 63.38 13.00 13.50
$w_{c}=0.025$ constant 75.63 69.00 62.88 64.00 12.00 13.50
$w_{c}=0.05$ constant 76.88 70.13 67.25 67.75 13.00 11.00
$w_{c}=0.1$ constant 77.50 70.50 71.75 74.63 12.00 11.50
$w_{c}=0.2$ constant 77.00 71.13 76.00 76.50 14.00 11.00
$w_{c}=0.4$ constant 76.13 70.50 81.00 80.88 14.00 10.75
$w_{c}=0.8$ constant 74.00 62.38 76.63 77.88 9.75 6.00
$w_{c}=1.6$ constant 74.38 66.63 82.25 82.13 21.75 14.00
$w_{c}=0.025$ cosine 76.88 70.75 81.50 79.50 17.25 11.50
$w_{c}=0.05$ cosine 74.25 64.13 81.63 80.63 21.00 11.75
$w_{c}=0.1$ cosine 76.88 72.13 82.00 81.75 22.00 13.50
$w_{c}=0.2$ cosine 74.00 67.13 83.00 82.00 25.25 11.25
$w_{c}=0.4$ cosine 77.75 70.38 82.00 82.00 22.50 17.25
$w_{c}=0.8$ cosine 75.63 67.25 82.75 81.63 21.75 14.75
$w_{c}=1.6$ cosine 68.88 66.00 83.63 82.50 18.25 9.75
Generative vs. representational alignment.
Increasing $w_{c}$ improves the linear separability of the final representations. Under the constant schedule, probe accuracy rises monotonically from $61.63\%$ at $w_{c}=0.0$ to $82.25\%$ at $w_{c}=1.6$. However, generative accuracy peaks early at $w_{c}=0.1$ ($77.50\%$) and degrades at higher values ($74.38\%$ at $w_{c}=1.6$), indicating that excessive contrastive pressure can disrupt the model’s generative capabilities.
Impact of loss counterpart scheduling on generalization.
The dual cosine schedule demonstrates superior efficiency in balancing these objectives. Even at low weights ($w_{c}=0.025$), the cosine schedule achieves high probe accuracy ($81.50\%$) comparable to aggressive constant schedules, while maintaining high generative performance.
Appendix O Example semantic concepts
@@
Appendix P Visualizing the Mechanism of PEFT via Attention Maps
@@
P.1 Distinct architectural signatures and tuning mechanisms
@@
Gemma3: overcoming representational degradation. As established in the main paper, Gemma3 models exhibit a high LSC but suffer from severe representational degradation, where the final-layer embeddings become poorly separated. The attention maps in Figure A.4 reveal the architectural reason and the solution. Gemma3 uses an efficient sliding window attention for the most part (narrow diagonal band) (layer 24 behaved differently for some reason).
@@
    With $\mathcal{L}_{\text{combined}}$ tuning, the contrastive loss induces a change that makes the model’s reasoning process visible. In the final layers, the attention map develops more structured, bright vertical stripes. Each well-defined stripe represents a targeted “read” operation, where the model globally accesses the compressed representation of visual features. This structured, cross-image comparison provides a clear mechanistic explanation for how the model overcomes its baseline shortcoming. The explicit contrastive pressure forces it to develop and execute this comparison strategy, repairing its degraded final representations and aligning with the recovery of final layer linear separability, as seen in Table 5.
@@
Figure A.4: Attention maps for Gemma3 4B. The contrastive loss in (b) visibly forces global cross-image attention in the final layers, a pattern not as pronounced in (a). This visualizes the mechanism that improves the model’s internal representations. The vertical axis represents the token attending from and the horizontal axis represents the token attending to.
Phi and Pixtral: representation refinement and feature enhancement. Phi and Pixtral employ a standard causal attention mask across all layers (Figures A.5 and A.6). Their path to success is not about fixing a shortcoming but refining existing processes.
With Phi possessing a higher-quality LSC from the outset, Phi’s fine-tuning is primarily a process of refinement and noise reduction. The attention maps visualize this as the patterns within each image block becoming more structured and less diffuse. Instead of a scattered focus across all visual patches, the model learns to consistently attend to the most salient features while ignoring irrelevant ones. This refined intra-image feature extraction allows for a cleaner signal to be passed to the final layers, enabling the model to better leverage its already-strong representations and surpass its LSC.
@@
Figure A.5: Attention maps for the Phi model after LoRA tuning with (a) the $\mathcal{L}_{\text{NT}}$ objective and (b) the $\mathcal{L}_{\text{combined}}$ objective. Both training methods result in a similar outcome: a refinement of attention over the model’s already high-quality initial representations. The patterns become visibly more structured and less diffuse, visualizing a mechanism of noise reduction and more targeted feature aggregation that enables the model to surpass its LSC. The vertical axis represents the token attending from and the horizontal axis represents the token attending to.
Pixtral’s success, even at baseline, stems from its inherent ability to perform vision representation refinement. The attention maps reveal the mechanism: an intensified focus on intra-image processing in later layers, visible as more structured diagonal blocks. Crucially, this visual signature is nearly identical when fine-tuning with either the standard $\mathcal{L}_{\text{NT}}$ or the $\mathcal{L}_{\text{combined}}$ objective. This equivalence demonstrates that standard next-token prediction is sufficient to fully engage Pixtral’s innate pathway for extracting and enhancing salient features.
@@
Figure A.6: Attention maps for the Pixtral model after LoRA tuning with (a) the $\mathcal{L}_{\text{NT}}$ objective and (b) the $\mathcal{L}_{\text{combined}}$ objective. The attention patterns are nearly identical, showing that both objectives activate the same underlying mechanism. The vertical axis represents the token attending from and the horizontal axis represents the token attending to.